# Time Series Forecasting with LLMs: Understanding and Enhancing Model Capabilities

### v3

@@ Abstract
Large language models (LLMs) have been applied in many fields and have developed rapidly in recent years. As a classic machine learning task, time series forecasting has recently been boosted by LLMs. Recent works treat large language models as zero-shot time series reasoners without further fine-tuning, which achieves remarkable performance. However, there are some unexplored research problems when applying LLMs for time series forecasting under the zero-shot setting. For instance, the LLMs’ preferences for the input time series are less understood. In this paper, by comparing LLMs with traditional time series forecasting models, we observe many interesting properties of LLMs in the context of time series forecasting. First, our study shows that LLMs perform well in predicting time series with clear patterns and trends but face challenges with datasets lacking periodicity. This observation can be explained by the ability of LLMs to recognize the underlying period within datasets, which is supported by our experiments. In addition, the input strategy is investigated and it is found that incorporating external knowledge and adopting natural language paraphrases substantially improve the predictive performance of LLMs for time series. Overall, our study contributes insight into LLMs’ advantages and limitations in time series forecasting under different conditions.
@@ 1 Introduction
Recently, large language models (LLMs) have been widely used and have achieved promising performance across various domains, such as health management, customer analysis, and text feature mining [16, 15, 12]. Time series forecasting requires extrapolation from sequential observations. Language models are designed to discern intricate concepts within temporally correlated sequences, and intuitively appear well-suited for this task. Hence, there exist some preliminary studies that apply LLMs to time series forecasting tasks [10, 17, 19].
However, currently the application of LLMs for time series forecasting is still in its early stage, and the boundaries of this research area are not yet well defined. There are many unexplored problems in this field. For example, existing research lacks exploration into how the performance of LLMs varies when faced with different types of time series inputs. This includes the effectiveness gap for LLMs in predicting data with seasonal and trending patterns versus data without such patterns.
To fill this research gap, in this paper, we focus on LLMs’ preferences for the input time series in time series forecasting under the zero shot prompting setting. Through experiments on both real and synthesized datasets, we find that LLMs perform better in time series with higher trend or seasonal strengths. Our observations also reveal that LLMs perform worse when there are multiple periods within datasets, which may be attributed to the fact that LLMs cannot capture the distinct periods within those datasets. To further discern the LLMs’ preferences for the specific segments of the input data, we design counterfactual experiments involving systematic permutations of input sequences. The findings suggest that LLMs are particularly sensitive to the segment of input sequences closest to the target output.
Based on the above findings, we want to further explore why LLMs forecast well on datasets with higher seasonal strengths. To this end, we require LLMs to tell the period of the datasets through multiple runs. We find that LLMs can mostly recognize the underlying period of a dataset. This can explain the findings of why large language models can forecast time series with high trends or seasonal intensities well, since they can obtain the seasonal pattern inside the datasets.
@@ 2.1 Large Language Model
Input Pre-processing. In this phase for time series forecasting with LLMs, we perform two pre-processing steps. First, numerical values are transformed into strings, a crucial step that significantly influences the model’s comprehension and data processing. For instance, a series like 0.123, 1.23, 12.3, 123.0 is reformatted to "1 2, 1 2 3, 1 2 3 0, 1 2 3 0 0", introducing spaces between digits and commas to delineate time steps, while decimal points are omitted to save token space. Second, tokenization is equally important, shaping the model’s pattern recognition capabilities. Unlike traditional methods such as byte-pair encoding (BPE) [13], which can disrupt numerical coherence, we use spacing digits which ensures individual tokenization, enhancing pattern discernment. Third, rescaling is employed to efficiently utilize tokens and manage large inputs by adjusting values so that a specific percentile aligns to 1. This facilitates the model’s exposure to varying digit counts and supports the generation of larger values, a testament to the nuanced yet critical nature of data preparation in leveraging LLMs for time series analysis.
@@ 2.2 Time Series Forecasting
Datasets. In this study, we primarily use Darts [11], a benchmark univariate dataset widely recognized in deep learning research, along with many baseline methods. Darts consists of eight real univariate time series datasets, including those with clear patterns, such as the AirPassengerDataset, and irregular datasets, such as the SunspotsDataset. Besides, we employ some other commonly used datasets, such as US Births Dataset[9], TSMC-Stock and Turkeypower datasets [10] and ETT [22] in Sections 5.1 and 5.2 to demonstrate the effectiveness of our proposed methods. A full description of those datasets can be seen in Appendix A.1.
@@ 3.1 Analyzing Method
We first compare the performance between LLMs and traditional time series forecasting methods, as shown in Table 9 and Table 10. It is shown that LLMs perform better within most datasets. GPT-4-turbo and Llama-2 perform relatively well on the AirPassengerdataset and the AusBeerdataset with low MAPE. Gemini outperforms GPT-3.5-turbo on time series forecasting and outperforms GPT-4-turbo on some datasets but is on par with GPT-4-turbo overall.
@@
where , and denote the trend component, the seasonal component and the residual component respectively. The presented indices indicate the trend’s strength and seasonality, providing a measure ranging up to 1. It is easy to find that a higher value indicates a stronger trend or seasonality within the time series. Throughout this paper, we use the word "higher strength" to represent the comparison of the strengths between different datasets. The assessment of strength is not based on a fixed level, as the concepts of "strong" and "weak" vary across different datasets and scenarios.
@@ 3.2 Preferences for Input Sequences
| GPT4-MAPE | 1.00000 | 0.987398 | -0.020637 | -0.681440 |
| GPT3.5-MAPE | 0.987398 | 1.00000 | -0.115087 | -0.669983 |
| Trend Strength | -0.020637 | -0.115087 | 1.00000 | 0.508980 |
| Seasonal Strength | -0.681440 | -0.669983 | 0.508980 | 1.00000 |
@@ 3.2.1 Implementation Details
Refer to subsubsection A.2.1 for detailed information.
Synthesized Datasets: To further validate our findings and investigate the influence of the number of periods on model performance, we generate a dataset using the function . ranges from 0 to 20 and follows the normal distribution . Refer to subsubsection A.2.1 for detailed information.
@@ 3.2.2 Key Findings
After computing the Pearson correlation coefficients (PCC), we observe a nearly strong correlation between the strengths and model performance, showing that LLMs perform better when the input time series has a higher trend and seasonal strength, which is shown in Table 1.In the context of multi-period time series, the model performance worsens as the number of periods increases. It indicates that LLMs may have difficulty recognizing the multiple periods inherent in such datasets. This potentially stems from their inability to adequately capture long-term periods, which can be supported in Figure 8. Besides, for counterfactual analysis, as shown in Figure 2 and Figure 3, there is a noticeable increase in MAPE values when Gaussian noise is added to the latter segments, while the perturbation of the first part of the sequence has little effect on the prediction performance. Our findings reveal that LLMs are more sensitive to the end of input time series when forecasting. We show our full results in Figure 6 and Figure 7. As we move to the right along the x-axis, the closer it gets to the output sequence. We defer the details in Appendix A.
It is also found that the initial part of the sequence has the least impact on the prediction accuracy. For the datasets with high seasonal strengths over 85%, such as WoolyDataset, and MonthlymilkDataset, more than 80% of the length of the time series has almost no effect on the model performance.
4 Why do LLMs Forecast Well on Data with Higher Seasonal Strengths?
@@
This approach is grounded in the hypothesis that LLMs’ proficiency in handling datasets with distinct seasonal attributes. By explicitly prompting LLMs to predict the dataset’s period, we aim to leverage their inherent ability to discern and extrapolate from complex patterns, which sheds light on the mechanisms that underpin their superior performance in such contexts.
@@ 4.1 Implementation Details
To explore the phenomenon that LLMs forecast well on datasets with higher seasonal strengths, we design experiments to verify this phenomenon. We tokenize the input sequence and let the LLMs output the period directly. We use GPT-3.5-turbo, GPT-4-turbo and Gemini-1.0-Pro to predict the periods. We have chosen five datasets with their seasonal strengths exceeding 85%. These datasets are readily available with clear seasonal patterns. In contrast, determining the specific periods of other irregular datasets is challenging, as they have no specific cycles. We record the predicted periods ten times and identify the mode period, which is the most frequently predicted value. We then compare the mode of these ten results with the real period. The mode is selected as the evaluation metric because, when considering the usage characteristics of LLMs, the output of this number best represents the model’s normal performance. We defer our details of the prompt in subsection A.3, and the results are shown in Table 2.

### v4

@@ Abstract
Large language models (LLMs) have been applied in many fields and have developed rapidly in recent years. As a classic machine learning task, time series forecasting has recently been boosted by LLMs. Recent works treat large language models as zero-shot time series reasoners without further fine-tuning, which achieves remarkable performance. However, some unexplored research problems exist when applying LLMs for time series forecasting under the zero-shot setting. For instance, the LLMs’ preferences for the input time series are less understood. In this paper, by comparing LLMs with traditional time series forecasting models, we observe many interesting properties of LLMs in the context of time series forecasting. First, our study shows that LLMs perform well in predicting time series with clear patterns and trends but face challenges with datasets lacking periodicity. This observation can be explained by the ability of LLMs to recognize the underlying period within datasets, which is supported by our experiments. In addition, the input strategy is investigated, and it is found that incorporating external knowledge and adopting natural language paraphrases substantially improve the predictive performance of LLMs for time series. Our study contributes insight into LLMs’ advantages and limitations in time series forecasting under different conditions.
@@ 1 Introduction
Recently, large language models (LLMs) have been widely used and have achieved promising performance across various domains, such as health management, customer analysis, and text feature mining [1, 2, 3, 4]. Time series forecasting requires extrapolation from sequential observations. Language models are designed to discern intricate concepts within temporally correlated sequences and intuitively appear well-suited for this task. Hence, some preliminary studies apply LLMs to time series forecasting tasks [5, 6, 7].
However, the application of LLMs for time series forecasting is still in its early stage, and the boundaries of this research area are not yet well defined. There are many unexplored problems in this field. For example, existing research lacks exploration into how the performance of LLMs varies when faced with different types of time series inputs. This includes the effectiveness gap for LLMs in predicting data with seasonal and trending patterns versus data without such patterns.
To fill this research gap, in this paper, we focus on LLMs’ preferences for the input time series in time series forecasting under the zero-shot prompting setting. Experiments on both real and synthesized datasets show that LLMs perform better in time series with higher trend or seasonal strengths. Our observations also reveal that LLMs perform worse when there are multiple periods within datasets, which may be attributed to the fact that LLMs cannot capture distinct periods within those datasets. To further discern the LLMs’ preferences for the specific input data segments, we design counterfactual experiments involving systematic permutations of input sequences. The findings suggest that LLMs are particularly sensitive to the segment of input sequences closest to the target output.
Based on the above findings, we want to explore why LLMs forecast well on datasets with higher seasonal strengths. To this end, we require LLMs to tell the period of the datasets through multiple runs. We find that LLMs can mostly recognize the underlying period of a dataset. This can explain the findings of why large language models can forecast time series with high trends or seasonal intensities well since they can obtain the seasonal pattern inside the datasets.
@@ 2.1 Large Language Model
Input Pre-processing. In this phase of time series forecasting with LLMs, we perform two pre-processing steps. First, numerical values are transformed into strings, a crucial step that significantly influences the model’s comprehension and data processing. For instance, a series like 0.123, 1.23, 12.3, 123.0 is reformatted to ”1 2, 1 2 3, 1 2 3 0, 1 2 3 0 0”, introducing spaces between digits and commas to delineate time steps, while decimal points are omitted to save token space. Second, tokenization is equally important, shaping the model’s pattern recognition capabilities. Unlike traditional methods such as byte-pair encoding (BPE) [8], which can disrupt numerical coherence, we use spacing digits which ensures individual tokenization, enhancing pattern discernment. Third, rescaling is employed to efficiently utilize tokens and manage large inputs by adjusting values so that a specific percentile aligns to 1. This facilitates the model’s exposure to varying digit counts and supports the generation of larger values, a testament to the nuanced yet critical nature of data preparation in leveraging LLMs for time series analysis.
@@ 2.2 Time Series Forecasting
Datasets. In this study, we primarily use Darts [10], a benchmark univariate dataset widely recognized in deep learning research, along with many baseline methods. Darts consists of eight real univariate time series datasets, including those with clear patterns, such as the AirPassengerDataset, and irregular datasets, such as the SunspotsDataset. Besides, we employ some other commonly used datasets, such as US Births Dataset[11], TSMC-Stock and Turkeypower datasets [5] and ETT [12] in Sections 5.2 and 5.3 to demonstrate the effectiveness of our proposed methods. A full description of those datasets can be seen in Section 5.1.
@@ 3.1 Analyzing Method
We first compare the performance between LLMs and traditional time series forecasting methods, as shown in Table 2. It is shown that LLMs perform better within most datasets. GPT-4-turbo and Llama-2 perform relatively well on the AirPassengerdataset and the AusBeerdataset with low MAPE. Gemini outperforms GPT-3.5-turbo on time series forecasting and outperforms GPT-4-turbo on some datasets but is on par with GPT-4-turbo overall.
@@
where , and denote the trend component, the seasonal component and the residual component respectively. The presented indices indicate the trend’s strength and seasonality, providing a measure ranging up to 1. It is easy to find that a higher value indicates a stronger trend or seasonality within the time series. Throughout this paper, we use the word ”higher strength” to represent the comparison of the strengths between different datasets. The assessment of strength is not based on a fixed level, as the concepts of ”strong” and ”weak” vary across different datasets and scenarios.
@@ 3.2 Preferences for Input Sequences
| GPT4-MAPE | ||||
| GPT3.5-MAPE | ||||
| Trend Strength | ||||
| Seasonal Strength |
@@ 3.2.1 Implementation Details
Synthesized Datasets: To further validate our findings and investigate the influence of the number of periods on model performance, we generate a dataset using the function . ranges from 0 to 20 and follows the normal distribution .
@@ 3.2.2 Key Findings
After computing the Pearson correlation coefficients (PCC), we observe a nearly strong correlation between the strengths and model performance, showing that LLMs perform better when the input time series has a higher trend and seasonal strength, which is shown in Table 1. In the context of multi-period time series, the model performance worsens as the number of periods increases. It indicates that LLMs may have difficulty recognizing the multiple periods inherent in such datasets. Besides, for counterfactual analysis, as shown in Figure 2 and Figure 3, there is a noticeable increase in MAPE values when Gaussian noise is added to the latter segments, while the perturbation of the first part of the sequence has little effect on the prediction performance. Our findings reveal that LLMs are more sensitive to the end of input time series when forecasting. We show our full results in Figure 2 and Figure 3. As we move to the right along the x-axis, the closer it gets to the output sequence. It is also found that the initial part of the sequence has the least impact on the prediction accuracy. For the datasets with high seasonal strengths over 85%, such as WoolyDataset, and MonthlymilkDataset, more than 80% of the length of the time series has almost no effect on the model performance.
4 Why do LLMs Forecast Well on Data with Higher
Seasonal Strengths?
@@ 4 Why do LLMs Forecast Well on Data with Higher
This approach is grounded in the hypothesis that LLMs are proficient in handling datasets with distinct seasonal attributes. By explicitly prompting LLMs to predict the dataset’s period, we aim to leverage their inherent ability to discern and extrapolate from complex patterns, which sheds light on the mechanisms that underpin their superior performance in such contexts.
| AirPassengers | AusBeer | ||
| Method | MSE / MAE / MAPE | Method | MSE / MAE / MAPE |
| Exponential Smoothing | 2007.67 / 37.91 / 8.10 | Exponential Smoothing | 703.26 / 22.80 / 5.44 |
| SARIMA | 2320.47 / 39.80 / 8.46 | SARIMA | 475.53 / 19.07 / 4.49 |
| Cyclical Regression | 2028.37 / 36.70 / 8.52 | Cyclical Regression | 989.31 / 26.29 / 6.13 |
| AutoARIMA | 8702.09 / 68.52 / 13.98 | AutoARIMA | 550.05 / 18.84 / 4.41 |
| FFT | 3274.46 / 46.38 / 10.59 | FFT | 7682.56 / 73.74 / 17.44 |
| StatsForecastAutoARIMA | 2952.52 / 45.41 / 9.71 | StatsForecastAutoARIMA | 559.46 / 20.56 / 4.86 |
| Naive Mean | 47703.65 / 204.25 / 44.61 | Naive Mean | 1885.72 / 30.66 / 6.68 |
| Naive Seasonal | 6032.80 / 62.87 / 14.18 | Naive Seasonal | 10828.02 / 96.35 / 23.39 |
| Naive Drift | 6505.79 / 72.21 / 17.50 | Naive Drift | 18507.61 / 128.23 / 30.91 |
| Naive Moving Average | 6032.80 / 62.87 / 14.18 | Naive Moving Average | 10828.02 / 96.35 / 23.39 |
| N-Beats | 3994.55 / 54.95 / 12.81 | N-Beats | 250.61 / 14.42 / 3.53 |
| DeepAR | 184222.64 / 421.99 / 98.42 | DeepAR | 16197.17 / 40.23 / 9.89 |
| Prophet | 7345.31 / 43.87 / 8.62 | Prophet | 6323.89 / 28.76 / 6.92 |
| LLMTime with GPT-3.5-Turbo | 6244.07 / 61.39 / 14.43 | LLMTime with GPT-3.5-Turbo | 841.68 / 23.59 / 5.62 |
| LLMTime with GPT-4-Turbo | 1317.9 / 55.49 / 11.18 | LLMTime with GPT-4-Turbo | 513.49 / 18.57 / 4.28 |
| LLMTime with Gemini-1.0-pro | 6392.21 / 63.57 / 14.03 | LLMTime with Gemini-1.0-pro | 397.78 / 14.36 / 3.27 |
| LLMtime with Llama-2 | 1286.25 / 28.04 / 6.07 | LLMtime with Llama-2 | 644.82 / 17.88 / 4.08 |
| MonthlyMilk | Sunspots | ||
| Method | MSE / MAE / MAPE | Method | MSE / MAE / MAPE |
| Exponential Smoothing | 564.94 / 20.23 / 2.41 | Exponential Smoothing | 326750.49 / 499.78 / 3129.63 |
| SARIMA | 1289.76 / 32.78 / 3.87 | SARIMA | 2902.72 / 45.75 / 466.99 |
| Cyclical Regression | 3631.53 / 56.15 / 6.60 | Cyclical Regression | 3917.76 / 47.84 / 274.31 |
| AutoARIMA | 2682.67 / 42.82 / 5.20 | AutoARIMA | 4695.67 / 58.47 / 709.23 |
| FFT | 3453.96 / 45.62 / 5.48 | FFT | 3784.56 / 49.81 / 150.32 |
| StatsForecastAutoARIMA | 186.14 / 10.64 / 1.28 | StatsForecastAutoARIMA | 8406.55 / 72.99 / 95.18 |
| Naive Mean | 19893.07 / 127.33 / 14.46 | Naive Mean | 4120.40 / 49.84 / 267.22 |
| Naive Seasonal | 4870.40 / 56.00 / 6.31 | Naive Seasonal | 4440.63 / 56.78 / 688.58 |
| Naive Drift | 3998.11 / 56.06 / 6.52 | Naive Drift | 5032.77 / 60.40 / 724.88 |
| Naive Moving Average | 4870.40 / 56.00 / 6.31 | Naive Moving Average | 4440.63 / 56.78 / 688.58 |
| N-Beats | 3140.89 / 51.57 / 6.07 | N-Beats | 4877.59 / 56.58 / 105.55 |
| DeepAR | 728289.50 / 851.30 / 99.22 | DeepAR | 3421.02 / 48.93 / 132.76 |
| Prophet | 663.41 / 25.76 / 2.92 | Prophet | 6303.57 / 76.83 / 67.97 |
| LLMTime with GPT-3.5-Turbo | 7507.13 / 66.28 / 112.77 | LLMTime with GPT-3.5-Turbo | 6556.55 / 58.95 / 217.94 |
| LLMTime with GPT-4-Turbo | 4442.18 / 50.75 / 172.82 | LLMTime with GPT-4-Turbo | 3374.70 / 41.87 / 321.11 |
| LLMTime with Gemini-1.0-pro | 628.98 / 17.01 / 1.99 | LLMTime with Gemini-1.0-pro | 626.03 / 14.94 / 1.73 |
| LLMtime with Llama-2 | 3410.20 / 41.40 / 240.25 | LLMtime with Llama-2 | 4467.67 / 48.95 / 91.79 |
| WineDataset | WoolyDataset | ||
| Method | MSE / MAE / MAPE | Method | MSE / MAE / MAPE |
| Exponential Smoothing | 23709576.52 / 3370.78 / 14.23 | Exponential Smoothing | 24925885.81 / 3548.19 / 14.98 |
| SARIMA | 1150166.94 / 966.57 / 20.76 | SARIMA | 812352.21 / 759.07 / 16.37 |
| Cyclical Regression | 7873785.27 / 2148.24 / 8.52 | Cyclical Regression | 1032574.82 / 962.72 / 22.14 |
| AutoARIMA | 698661.90 / 646.03 / 14.07 | AutoARIMA | 838852.91 / 786.25 / 16.84 |
| FFT | 1031170.45 / 867.83 / 18.60 | FFT | 1012255.35 / 945.20 / 20.80 |
| StatsForecastAutoARIMA | 20040877.37 / 2853.17 / 12.05 | StatsForecastAutoARIMA | 917617.19 / 858.57 / 18.91 |
| Naive Mean | 11557786.19 / 2200.04 / 8.80 | Naive Mean | 816762.31 / 764.73 / 16.12 |
| Naive Seasonal | 879447.22 / 724.23 / 15.52 | Naive Seasonal | 1051110.81 / 982.25 / 22.19 |
| Naive Drift | 9609576.04 / 1833.38 / 7.36 | Naive Drift | 812352.21 / 759.07 / 16.37 |
| Naive Moving Average | 9070696.99 / 1719.17 / 6.90 | Naive Moving Average | 1032574.82 / 962.72 / 22.14 |
| N-Beats | 5418377.00 / 1887.30 / 7.68 | N-Beats | 653104.31 / 743.54 / 15.96 |
| DeepAR | 715027008.00 / 26236.14 / 89.91 | DeepAR | 243831.14 / 4897.85 / 94.89 |
| Prophet | 4846922.27 / 2201.57 / 8.27 | Prophet | 365241.98 / 891.70 / 34.65 |
| LLMTime with GPT-3.5-Turbo | 30488.60 / 388.28 / 15.83 | LLMTime with GPT-3.5-Turbo | 526903.08 / 574.58 / 12.00 |
| LLMTime with GPT-4-Turbo | 22488.17 / 253.08 / 9.98 | LLMTime with GPT-4-Turbo | 942987.19 / 871.64 / 18.55 |
| LLMTime with Gemini-1.0-pro | 258584.78 / 3645.23 / 14.60 | LLMTime with Gemini-1.0-pro | 64.92 / 6.39 / 7.04 |
| LLMtime with Llama-2 | 951194.94 / 240.08 / 9.45 | LLMtime with Llama-2 | 675062.52 / 736.04 / 15.83 |
| HeartRateDataset | Weather | ||
| Method | MSE / MAE / MAPE | Method | MSE / MAE / MAPE |
| Exponential Smoothing | 11.16 / 1.38 / 1.49 | Exponential Smoothing | 1684.38 / 31.60 / 6.79 |
| SARIMA | 12.98 / 1.34 / 1.61 | SARIMA | 1943.81 / 33.33 / 7.09 |
| Cyclical Regression | 13.58 / 1.31 / 1.20 | Cyclical Regression | 1700.73 / 30.77 / 7.15 |
| AutoARIMA | 13.26 / 1.25 / 1.39 | AutoARIMA | 7315.10 / 57.44 / 11.70 |
| FFT | 13.95 / 1.16 / 1.34 | FFT | 2752.02 / 38.90 / 8.87 |
| StatsForecastAutoARIMA | 10.53 / 1.27 / 1.39 | StatsForecastAutoARIMA | 2479.55 / 38.06 / 8.16 |
| Naive Mean | 12.02 / 1.27 / 1.26 | Naive Mean | 39879.84 / 168.27 / 36.44 |
| Naive Seasonal | 10.55 / 1.32 / 1.31 | Naive Seasonal | 5057.47 / 52.81 / 11.89 |
| Naive Drift | 10.60 / 1.15 / 1.30 | Naive Drift | 5466.23 / 60.58 / 14.70 |
| Naive Moving Average | 12.13 / 1.27 / 1.34 | Naive Moving Average | 5057.47 / 52.81 / 11.89 |
| N-Beats | 72.11 / 7.10 / 7.40 | N-Beats | 4532.84 / 39.21 / 23.49 |
| DeepAR | 286.82 / 15.67 / 16.36 | DeepAR | 6325.75 / 35.97 / 16.59 |
| Prophet | 88.93 / 10.97 / 6.54 | Prophet | 3768.15 / 29.36 / 24.01 |
| LLMTime with GPT-3.5-Turbo | 76.83 / 7.15 / 7.42 | LLMTime with GPT-3.5-Turbo | 224.54 / 3.07 / 0.83 |
| LLMTime with GPT-4-Turbo | 988.14 / 26.57 / 29.22 | LLMTime with GPT-4-Turbo | 111.65 / 2.40 / 0.64 |
| LLMTime with Gemini-1.0-pro | 57.96 / 6.01 / 6.66 | LLMTime with Gemini-1.0-pro | 176.32 / 3.72 / 0.75 |
| LLMtime with Llama-2 | 75.58 / 7.11 / 7.94 | LLMtime with Llama-2 | 215.39 / 4.07 / 1.31 |
@@ 4.1 Implementation Details
To explore the phenomenon that LLMs forecast well on datasets with higher seasonal strengths, we design experiments to verify this phenomenon. We tokenize the input sequence and let the LLMs output the period directly. We use GPT-3.5-turbo, GPT-4-turbo, and Gemini-1.0-Pro to predict the periods. We have chosen five datasets with their seasonal strengths exceeding 85%. These datasets are readily available with clear seasonal patterns. In contrast, determining the specific periods of other irregular datasets is challenging, as they have no specific cycles. We record the predicted periods ten times and identify the mode period, which is the most frequently predicted value. We then compare the mode of these ten results with the real period. The mode is selected as the evaluation metric because, when considering the usage characteristics of LLMs, the output of this number best represents the model’s normal performance. The results are shown in Table 3.